There are 3 fewer dimes than nickels. The collection totals $1.35. How many dimes are in the collection?
Answer by mananth(16946)

let nickels be x
dimes = x-3
..
value of coins = 135 cents
..
5x+10(x-3)=135
5x+10x-30=135
15x-30=135
add 30
15x-30+30=135+30
15x=165
divide by 15
15x/15=165/15
x=11 nickels
dimes will be x-3= 11-3=8
..
CHECK
11*5+10*8=135